OpenStudy (anonymous):

what is the exact value of tan15

OpenStudy (anonymous):

tan(A−B) = (tan A − tan B) / (1 + tan A tan B)

OpenStudy (anonymous):

Tan(45 - 15)

OpenStudy (anonymous):

OOPS Tan(45 - 30)

OpenStudy (anonymous):

(tan 45 - tan 30)/(1 + tan45*tan30)

OpenStudy (anonymous):

Tan 45 = 1 tan 30 = sqrt3/3

OpenStudy (anonymous):

(1 - sqrt3/3)(1 + sqrt3/3)

OpenStudy (anonymous):

how do you simplify that? the answer is supposed to be 2 - sqrt3

OpenStudy (anonymous):

Just a minute... I don't get that answer.. Let me go back thru my math

OpenStudy (anonymous):

i got up to where you got before too

OpenStudy (anonymous):

OK (1 = sqrt3/3) / (1 + sqrt3/3)

OpenStudy (anonymous):

Now you multiply by the conjugate which is (1 - sqrt3/3)/(1-sqrt3/3)

OpenStudy (anonymous):

(1 - sqrt3/3)(1 - sqrt3/3) --------------------- (1 + sqrt3/3)(1 - sqrt3/3)

OpenStudy (anonymous):

1 - sqrt3/3 - sqrt3/3 + 3/9 ------------------------ 1 - 3/9

OpenStudy (anonymous):

1 - 2sqrt3/3 + 1/3 ----------------- 2/3

OpenStudy (anonymous):

4/3 - 2sqrt3/3 ------------- 2/3

OpenStudy (anonymous):

(4 - 2sqrt3)/3 divided by 2/3 (4 - 2sqrt3)/3 times 3/2

OpenStudy (anonymous):

The threes would cancel (4 - 2sqrt3)/2

OpenStudy (anonymous):

divide the 2 into each term 2 - sqrt3

OpenStudy (anonymous):

how do you find the conjugate again?

OpenStudy (anonymous):

The conjugate is basically the denominator but take the opposite sign ex. denominator = 1 - sqrt5 conjugate 1+sqrt5/1+sqrt5

OpenStudy (anonymous):

perfect, thanks!

OpenStudy (anonymous):

UR welcome

OpenStudy (anonymous):

hey, you could use this http://web2.0calc.com/